# The inverse Kalman filter

### v4

@@ 2 Inverse Kalman filter
In this section, we introduce an exact algorithm for computing $\bm{\Sigma}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$, where $\mathbf{u}$ is an $N$-dimensional real-valued vector and $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]$ is an $N\times N$ covariance matrix of observations induced by a DLM with $q$ latent states in Eqs. (1)-(2). This algorithm is applicable to any DLM specified in Eqs. (1)-(2). Denote $\bm{\Sigma}=\mathbf{L}\mathbf{L}^{T}$, where the Cholesky factor $\mathbf{L}$ does not need to be explicitly computed in our approach. We develop fast algorithms to compute $\mathbf{\tilde{x}}=\mathbf{L}^{T}\mathbf{u}$ and $\mathbf{x}=\mathbf{L}\mathbf{\tilde{x}}$ by Lemma 1 and Lemma 2, respectively, each with $\mathcal{O}(q^{3}N)$ operations. Detailed proofs of these lemmas are available in the Supplementary Material. In the following, $u_{t}$, $x_{t}$, $\tilde{x}_{t}$ denote the $t$th element of the vector of $\mathbf{u}$, $\mathbf{x}$, and $\mathbf{\tilde{x}}$, respectively, and $L_{t^{\prime},t}$ denotes the $(t^{\prime},t)$th element of $\mathbf{L}$ for $t$ and $t^{\prime}=1,2,\dots,N$.
@@
with $\mathbf{B}_{t}=\mbox{cov}[\bm{\theta}_{t}\mid{\mathbf{y}}_{1:t-1}]$ and $Q_{t}=\mbox{var}[y_{t}\mid{\mathbf{y}}_{1:t-1}]$ from the Kalman filter in (26) and (27) in Appendix A, respectively. Then $\tilde{\mathbf{x}}=\mathbf{L}^{T}{\mathbf{u}}$.
@@
where $Q_{t}$ and $\mathbf{K}_{t}$ are from (27) and (6), respectively. Then $\mathbf{x}=\mathbf{L}\tilde{\mathbf{x}}$.
@@
While Algorithm 1 can be applied to DLM regardless of the noise variance, we do not recommend directly computing $\bm{\Sigma}\mathbf{u}$ in scenarios when the noise variance $V_{t}$ is zero or close to zero in Eq. (1). This is because when $V_{t}=0$, $Q_{t}$ in Eq. (27), the variance of the one-step-ahead predictive distribution, could be extremely small, which leads to large numerical errors in Eq. (28) due to unstable inversion of $Q_{t}$. To ensure robustness, we suggest computing $\mathbf{x}=(\bm{\Sigma}+V\mathbf{I}_{N}){\mathbf{u}}$ with a positive scalar $V$ and outputting $\mathbf{x}-V{\mathbf{u}}$ as the result of $\bm{\Sigma}{\mathbf{u}}$. Here, $\bm{\Sigma}+V\mathbf{I}_{N}$ can be interpreted as the covariance matrix of a DLM with noise variance $V_{t}=V$. Essentially, we introduce artificial noise to ensure that $Q_{t}$ computed in KF is at least $V$ for numerical stability. The result remains exact and independent of the choice of $V$.
@@
with $\prod^{t^{\prime}}_{l=t+1}\mathbf{G}_{l}=\mathbf{G}_{t^{\prime}}\mathbf{G}_{t^{\prime}-1}\dots\mathbf{G}_{t+1}$, $Q_{t}$ and $\mathbf{K}_{t}$ defined in (27) and (6), respectively.
@@ 3 Nonparametric estimation of particle interaction functions
An illustrative example of this framework is an unnormalized Vicsek model. Unlike the original Vicsek model [36], which enforces a constant speed for all particles, we allow velocity magnitudes to vary over time. Time-varying velocities are more sensible in many real-world scenarios, including cellular movements studied in Section 4.4. When the system approaches a high degree of confluence, the average velocity of the cells typically decreases because of cell proliferation. The 2-dimensional velocity of the $i^{\prime}$th particle at time $\tau$, $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, is updated by averaging the velocity of its neighbors from the previous time step. Specifically, for each direction $l=1,2$,
@@
The unnormalized Vicsek model in Eq. (14) is a special case of the general framework in Eq. (13) with scalar output $\mathbf{y}_{i}=v_{i^{\prime},l}(\tau)$, i.e. $D_{y}=1$, and the index of the $i$th observation $i=(\tau-1)n_{p}+i^{\prime}$ is a function of time $\tau$ and particle $i^{\prime}$. This model contains a single interaction, i.e. $J=1$, with a linear interaction function $z(d)=d$, where $d$ is the velocity component of the neighboring particle, and the factor loading being $1/p_{i^{\prime}}(\tau-1)$.
Numerous other physical models of self-propelled particles can also be formulated as in Eq. (13) with a parametric form of a particle interaction function [1]. Nonparametric estimation of particle interactions is preferred when the physical mechanism is unknown [18, 24], but the high computational cost limits its applicability to systems with large numbers of particles over long time periods.
@@
with the predictive mean and variance given by
@@
where $\bm{\Sigma}_{y}=\sum^{J}_{j=1}\mathbf{A}_{j}\bm{\Sigma}^{(u)}_{j}\mathbf{A}^{T}_{j}+\sigma^{2}_{0}\mathbf{I}_{\tilde{N}}$, $\bm{\Sigma}^{(u)}_{j}(d^{*})=[c_{j}(d^{(u)}_{1,j},d^{*}),\dots,c_{j}(d^{(u)}_{N_{j},j},d^{*})]^{T}$, and $\mathbf{r}$ represents additional system parameters.
@@ 4.2 Unnormalized Vicsek model
We first consider the unnormalized Vicsek model introduced in Section 3. For each particle $i^{\prime}$ at time $\tau$, after obtaining its velocity using Eq. (14), its position is updated as:
$\mathbf{s}_{i^{\prime}}(\tau)$ $=\mathbf{s}_{i^{\prime}}(\tau-1)+\mathbf{v}_{i^{\prime}}(\tau)h,\quad i=1^{\prime},\dots,n_{p},$ (23)
@@
where $h$ is the time step size. Particles are initialized with velocity $[v\cos(\phi_{i^{\prime}}(0)),v\sin(\phi_{i^{\prime}}(0))]^{T}$, where $\phi_{i^{\prime}}(0)$ is drawn from $\mbox{Unif}[-\pi,\pi]$ and $v={2}^{1/2}/2\approx 0.71$. Initial particle positions $\mathbf{s}_{i^{\prime}}(0)$ are uniformly sampled from $[0,{n_{p}}^{1/2}]\times[0,{n_{p}}^{1/2}]$ to keep consistent particle density across experiments. The goal is to estimate the interaction function nonparametrically, without assuming linearity. The interaction radius $r=0.5$ is estimated alongside other parameters, with results detailed in the Supplementary Material.
@@
Figure 4: Boxplots of NRMSE (Eq. (20)) for estimating the latent interaction function in the unnormalized Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ based on 20 experiments. Panel (a) uses the Matérn covariance with roughness parameter 2.5, and panel (b) uses the exponential covariance.
@@
We first compare the computational cost and accuracy of IKF-CG, CG, and direct computation for calculating the predictive mean in Eq. (18) and the marginal likelihood in Eq. (16) using the covariance in Eq. (15). Simulations are conducted with $n_{p}=100$ particles and noise variance $\sigma^{2}_{0}=0.2^{2}$. Panel (a) of Fig. 3 shows that the IKF-CG approach substantially outperforms both direct computation and the CG algorithm in computational time when predicting $n^{*}=100$ test inputs with varying time lengths $n_{\tau}$, ranging from 4 to 200. Figure 3(b) compares the NRMSE of the predictive mean between the direct computation and IKF-CG method when the number of observations ranges from 400 to 2,000, corresponding to $n_{\tau}$ from 4 to 20. The two methods yield nearly identical predictive errors, with negligible differences in NRMSE. Figure 3(c) shows a comparison of the log-likelihood values computed via direct computation and IKF-CG, where the latter employs the log-determinant approximation detailed in the Supplementary Material. Both methods produce similar log-likelihood values across different sample sizes.
@@
Figure 5: Uncertainty assessment of predicted interaction function in the unnormalized Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ using (a) Matérn covariance function in Eq. (15) and (b) exponential covariance function. Bars represent average lengths of 95% posterior credible intervals (Eq. (21)). Dots indicate the proportions covered in the 95% intervals (Eq. (22)), with optimal coverage (0.95) shown as the purple dashed line.
@@
Figure 6: (a) Predicted (blue dashed lines) and true (pink solid lines) interaction functions with $95\%$ posterior credible interval (grey shaded area) for the unnormalized Vicsek model with $n_{\tau}=900$, $T=10$, and $\sigma_{0}^{2}=0.1^{2}$, using Matérn covariance function with $\nu=2.5$. (b) Trajectories of $45$ sample particles over $50$ time frames using estimated (blue dashed lines) and true (pink solid lines) interaction functions with identical initial positions and noise samples.
@@
Panels (a) and (b) of Figs. 4-5 show the predictive performance of particle interactions using the Matérn covariance with roughness parameter $\nu=2.5$ and the exponential covariance (Matérn with $\nu=0.5$), respectively, for noise variance $\sigma_{0}^{2}=0.1^{2}$. Results for $\sigma_{0}^{2}=0.2^{2}$ are provided in the Supplementary Material. Across all scenarios, NRMSEs remain low, with improved accuracy for larger datasets and smaller noise variances. The decrease in the average length of the $95\%$ posterior credible interval with increasing sample size, along with the relatively small interval span compared to the output range, indicates improved prediction confidence with more observations. Moreover, the coverage proportion for the $95\%$ posterior credible interval is close to the nominal $95\%$ level in all cases, validating the reliability of the uncertainty quantification. By comparing panels (a) and (b) in Figs. 4-5, we find that the Matérn covariance in Eq. (15) yields lower NRMSE and narrower $95\%$ posterior credible intervals than the exponential kernel. This improvement is due to the smoother latent process induced by Eq. (15), which is twice mean-squared differentiable, while the process with an exponential kernel is not differentiable.
@@ 4.3 A modified Vicsek model with multiple interactions
Various extensions of the Vicsek model have been studied to capture more complex particle dynamics [1, 9]. For illustration, we consider a modified Vicsek model with two interaction functions. The 2-dimensional velocity $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, corresponding to the output in Eq. (13), is updated according to:
@@
We simulate 12 scenarios, each replicated $E=20$ times, using the same number of particles, time frame, and noise level as in the unnormalized Vicsek model in Section 4.2. The predictive performance of the IKF-CG is evaluated using $n^{*}=200$ test inputs evenly spaced across the input domain of each interaction function. We focus on the model with the Matérn kernel in Eq. (15), and results for the exponential kernel are detailed in the Supplementary Material. Consistent with the results of the unnormalized Vicsek model, we find models with the Matérn kernel in Eq. (15) are more accurate due to the smooth prior imposed on the interaction function.
@@ 4.4 Estimating cell-cell interactions on anisotropic substrates
where $n_{p}(\tau)$ is the cell count at time $\tau$, $l=1,2$ correspond to horizontal and vertical directions, respectively, and $\epsilon_{i^{\prime},l}(\tau)\sim\mathcal{N}(0,\sigma^{2}_{0,l}(\tau))$ denotes the noise. Inspired by the modified Vicsek model, we set $d_{k,l}=v_{k,l}(\tau-1)$. To account for velocity decay caused by increasing cell confluence, we model the noise variances as $\sigma^{2}_{0,l}(\tau)=\omega_{l}\sigma^{2}_{v,l}(\tau-1)$, where $\sigma^{2}_{v,l}(\tau-1)$ is the sample velocity variance at time $(\tau-1)$, and $\omega_{l}$ is a parameter estimated from data for $l=1,2$. The neighboring set $ne_{i^{\prime},l}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r_{l}\mbox{ and }\mathbf{v}_{i^{\prime}}(\tau-1)^{T}\mathbf{v}_{k}(\tau-1)>0\}$, where $r_{l}$ denotes the interaction radius of direction $l$ for $l=1,2$, excludes particles moving in opposite directions, reflecting the observed gliding and intercalating behavior of cells [37].
@@
Table 2: One-step-ahead prediction performance on the testing dataset. Here $\mbox{RMSE}=\{\sum_{\tau=1}^{n_{\tau}^{*}}\sum_{i^{\prime}=1}^{n_{p}(\tau)}(\hat{v}_{i^{\prime},l}(\tau)-v_{i^{\prime},l}(\tau))^{2}/\sum_{\tau=1}^{n_{\tau}^{*}}n_{p}(\tau)\}^{1/2}$, with $n_{\tau}^{*}$ denoting the number of testing time frames. L(95%) and P(95%) are computed similarly to those in Eqs. (21) - (22), but with the predictive interval of $z$ replaced by that of $y$, where $y$ represents the velocity in each direction.

### v5

@@ 2 Inverse Kalman filter
In this section, we introduce an exact algorithm for computing $\bm{\Sigma}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$, where $\mathbf{u}$ is an $N$-dimensional real-valued vector and $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]$ is an $N\times N$ covariance matrix of observations induced by a DLM with $q$ latent states in (1)-(2). This algorithm is applicable to any DLM specified in (1)-(2). Denote $\bm{\Sigma}=\mathbf{L}\mathbf{L}^{T}$, where the Cholesky factor $\mathbf{L}$ does not need to be explicitly computed in our approach. We develop fast algorithms to compute $\mathbf{\tilde{x}}=\mathbf{L}^{T}\mathbf{u}$ and $\mathbf{x}=\mathbf{L}\mathbf{\tilde{x}}$ by Lemma 1 and Lemma 2, respectively, each with $\mathcal{O}(q^{3}N)$ operations. Detailed proofs of these lemmas are available in the Supplementary Material. In the following, $u_{t}$, $x_{t}$, $\tilde{x}_{t}$ denote the $t$th element of the vector of $\mathbf{u}$, $\mathbf{x}$, and $\mathbf{\tilde{x}}$, respectively, and $L_{t^{\prime},t}$ denotes the $(t^{\prime},t)$th element of $\mathbf{L}$ for $t$ and $t^{\prime}=1,2,\dots,N$.
@@
with $\mathbf{B}_{t}=\mbox{cov}[\bm{\theta}_{t}\mid{\mathbf{y}}_{1:t-1}]$ and $Q_{t}=\mbox{var}[y_{t}\mid{\mathbf{y}}_{1:t-1}]$ from the Kalman filter in (A1) and (A2) in Appendix A, respectively. Then $\tilde{\mathbf{x}}=\mathbf{L}^{T}{\mathbf{u}}$.
@@
where $Q_{t}$ and $\mathbf{K}_{t}$ are from (A2) and (6), respectively. Then $\mathbf{x}=\mathbf{L}\tilde{\mathbf{x}}$.
@@
While Algorithm 1 can be applied to DLM regardless of the noise variance, we do not recommend directly computing $\bm{\Sigma}\mathbf{u}$ in scenarios when the noise variance $V_{t}$ is zero or close to zero in (1). This is because when $V_{t}=0$, $Q_{t}$ in (A2), the variance of the one-step-ahead predictive distribution, could be extremely small, which leads to large numerical errors in (A3) due to unstable inversion of $Q_{t}$. To ensure robustness, we suggest computing $\mathbf{x}=(\bm{\Sigma}+V\mathbf{I}_{N}){\mathbf{u}}$ with a positive scalar $V$ and outputting $\mathbf{x}-V{\mathbf{u}}$ as the result of $\bm{\Sigma}{\mathbf{u}}$. Here, $\bm{\Sigma}+V\mathbf{I}_{N}$ can be interpreted as the covariance matrix of a DLM with noise variance $V_{t}=V$. Essentially, we introduce artificial noise to ensure that $Q_{t}$ computed in KF is at least $V$ for numerical stability. The result remains exact and independent of the choice of $V$.
@@
with $\prod^{t^{\prime}}_{l=t+1}\mathbf{G}_{l}=\mathbf{G}_{t^{\prime}}\mathbf{G}_{t^{\prime}-1}\dots\mathbf{G}_{t+1}$, $Q_{t}$ and $\mathbf{K}_{t}$ defined in (A2) and (6), respectively.
@@ 3 Nonparametric estimation of particle interaction functions
An illustrative example of this framework is an unnormalized Vicsek model. Unlike the original Vicsek model [36], which enforces a constant speed for all particles, we allow velocity magnitudes to vary over time. Time-varying velocities are more sensible in many real-world scenarios, including cellular movements studied in Section 4.4. When the system approaches a high degree of confluence, the average velocity of the cells typically decreases because of cell proliferation. Assume the 2-dimensional velocity of the $i^{\prime}$th particle at time $\tau$, $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, is updated by averaging the velocity of its neighbors from the previous time step. Specifically, for each direction $l=1,2$,
@@
The unnormalized Vicsek model in (14) is a special case of the general framework in (13) with scalar output $\mathbf{y}_{i}=v_{i^{\prime},l}(\tau)$, i.e. $D_{y}=1$, and the index of the $i$th observation $i=2(\tau-1)n_{p}+2(i^{\prime}-1)+l$ is a function of time $\tau$ and particle $i^{\prime}$. This model contains a single interaction, i.e. $J=1$, with a linear interaction function $z(d)=d$, where $d$ is the velocity component of the neighboring particle, and the factor loading being $1/p_{i^{\prime}}(\tau-1)$.
Numerous other physical models of self-propelled particles can also be formulated as in (13) with a parametric form of a particle interaction function [1]. Nonparametric estimation of particle interactions is preferred when the physical mechanism is unknown [18, 24], but the high computational cost limits its applicability to systems with large numbers of particles over long time periods.
@@
where the predictive mean and variance are given by
@@
with $\bm{\Sigma}_{y}=\sum^{J}_{j=1}\mathbf{A}_{j}\bm{\Sigma}^{(u)}_{j}\mathbf{A}^{T}_{j}+\sigma^{2}_{0}\mathbf{I}_{\tilde{N}}$, $\bm{\Sigma}^{(u)}_{j}(d^{*})=[c_{j}(d^{(u)}_{1,j},d^{*}),\dots,c_{j}(d^{(u)}_{N_{j},j},d^{*})]^{T}$, and $\mathbf{r}$ representing additional system parameters.
@@ 4.2 Unnormalized Vicsek model
We first consider the unnormalized Vicsek model introduced in Section 3. For each particle $i^{\prime}$ at time $\tau$, after obtaining its velocity using (14), its position is updated as:
$\mathbf{s}_{i^{\prime}}(\tau)$ $=\mathbf{s}_{i^{\prime}}(\tau-1)+\mathbf{v}_{i^{\prime}}(\tau)h,\quad i^{\prime}=1,\dots,n_{p},$ (23)
@@
where $h$ is the time step size. Particles are initialized with velocity $[v\cos(\phi_{i^{\prime}}(0)),v\sin(\phi_{i^{\prime}}(0))]^{T}$, where $\phi_{i^{\prime}}(0)$ is drawn from $\mbox{Unif}[-\pi,\pi]$ and $v={2}^{1/2}/2\approx 0.71$. Initial particle positions $\mathbf{s}_{i^{\prime}}(0)$ in (23) are uniformly sampled from $[0,{n_{p}}^{1/2}]\times[0,{n_{p}}^{1/2}]$ to keep consistent particle density across experiments. The goal is to estimate the interaction function nonparametrically, without assuming linearity. The interaction radius $r=0.5$ is estimated alongside other parameters, with the estimation results detailed in the Supplementary Material.
@@
Figure 4: Boxplots of NRMSE (20) for estimating the latent interaction function in the unnormalized Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ based on 20 experiments. Panel (a) uses the Matérn covariance with roughness parameter 2.5, and panel (b) uses the exponential covariance.
@@
We first compare the computational cost and accuracy of IKF-CG, CG, and direct computation for calculating the predictive mean in (18) and the marginal likelihood in (16) using the covariance in (15). Simulations are conducted with $n_{p}=100$ particles and noise variance $\sigma^{2}_{0}=0.2^{2}$. Panel (a) of Fig. 3 shows that the IKF-CG approach substantially outperforms both direct computation and the CG algorithm in computational time when predicting $n^{*}=100$ test inputs with varying time lengths $n_{\tau}$, ranging from 4 to 200. Figure 3(b) compares the NRMSE of the predictive mean between the direct computation and IKF-CG method when the number of observations ranges from 400 to 2,000, corresponding to $n_{\tau}$ from 4 to 20. The two methods yield nearly identical predictive errors, with negligible differences in NRMSE. Figure 3(c) shows a comparison of the log-likelihood values computed via direct computation and IKF-CG, where the latter employs the log-determinant approximation detailed in the Supplementary Material. Both methods produce similar log-likelihood values across different sample sizes.
@@
Figure 5: Uncertainty assessment of predicted interaction function in the unnormalized Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ using (a) Matérn covariance function in (15) and (b) exponential covariance function. Bars represent average lengths of 95% posterior credible intervals (21). Dots indicate the proportions covered in the 95% intervals (22), with the optimal coverage (0.95) shown as the purple dashed lines.
@@
Figure 6: (a) Predicted (blue dashed lines) and true (pink solid lines) interaction functions with $95\%$ posterior credible interval (grey shaded area) for the unnormalized Vicsek model with $n_{\tau}=900$, $T=10$, and $\sigma_{0}^{2}=0.1^{2}$, using Matérn covariance function with $\nu=2.5$. (b) Trajectories of $45$ randomly sampled particles over $50$ time frames using estimated (blue dashed lines) and true (pink solid lines) interaction functions with identical initial positions and noise samples.
@@
Panels (a) and (b) of Figs. 4-5 show the predictive performance of particle interactions using the Matérn covariance with roughness parameter $\nu=2.5$ and the exponential covariance (Matérn with $\nu=0.5$), respectively, for noise variance $\sigma_{0}^{2}=0.1^{2}$. Results for $\sigma_{0}^{2}=0.2^{2}$ are provided in the Supplementary Material. Across all scenarios, NRMSEs remain low, with improved accuracy for larger datasets and smaller noise variances. The decrease in the average length of the $95\%$ posterior credible interval with increasing sample size, along with the relatively small interval span compared to the output range, indicates improved prediction confidence with more observations. Moreover, the coverage proportion for the $95\%$ posterior credible interval is close to the nominal $95\%$ level in all cases, validating the reliability of the uncertainty quantification. By comparing panels (a) and (b) in Figs. 4-5, we find that the Matérn covariance in (15) yields lower NRMSE and narrower $95\%$ posterior credible intervals than the exponential kernel. This improvement is due to the smoother latent process induced by (15), which is twice mean-squared differentiable, while the process with an exponential kernel is not differentiable.
@@ 4.3 A modified Vicsek model with multiple interactions
Various extensions of the Vicsek model have been studied to capture more complex particle dynamics [1, 9]. For illustration, we consider a modified Vicsek model with two interaction functions. The 2-dimensional velocity $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, corresponding to the output in (13), is updated according to:
@@
We simulate 12 scenarios, each replicated $E=20$ times, using the same number of particles, time frame, and noise level as in the unnormalized Vicsek model in Section 4.2. The predictive performance of the IKF-CG is evaluated using $n^{*}=200$ test inputs evenly spaced across the input domain of each interaction function. We focus on the model with the Matérn kernel in (15), and results for the exponential kernel are detailed in the Supplementary Material. Consistent with the results of the unnormalized Vicsek model, we find models with the Matérn kernel in (15) are more accurate due to the smooth prior imposed on the interaction function.
@@ 4.4 Estimating cell-cell interactions on anisotropic substrates
where $n_{p}(\tau)$ is the cell count at time $\tau$, $l=1,2$ correspond to horizontal and vertical directions, respectively, and $\epsilon_{i^{\prime},l}(\tau)\sim\mathcal{N}(0,\sigma^{2}_{0,l}(\tau))$ denotes the noise. Inspired by the modified Vicsek model, we set $d_{k,l}=v_{k,l}(\tau-1)$. To account for velocity decay caused by increasing cell confluence, we model the noise variances as $\sigma^{2}_{0,l}(\tau)=\omega_{l}\sigma^{2}_{v,l}(\tau-1)$, where $\sigma^{2}_{v,l}(\tau-1)$ is the sample velocity variance at time $(\tau-1)$, and $\omega_{l}$ is a parameter estimated from data for $l=1,2$. The neighboring set in (25) is formed as $ne_{i^{\prime},l}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r_{l}\mbox{ and }\mathbf{v}_{i^{\prime}}(\tau-1)^{T}\mathbf{v}_{k}(\tau-1)>0\}$, where $r_{l}$ denotes the interaction radius of direction $l$ for $l=1,2$, excludes particles moving in opposite directions, reflecting the observed gliding and intercalating behavior of cells [37].
@@
Table 2: One-step-ahead prediction performance on the testing dataset. Here $\mbox{RMSE}=\{\sum_{\tau=1}^{n_{\tau}^{*}}\sum_{i^{\prime}=1}^{n_{p}(\tau)}(\hat{v}_{i^{\prime},l}(\tau)-v_{i^{\prime},l}(\tau))^{2}/\sum_{\tau=1}^{n_{\tau}^{*}}n_{p}(\tau)\}^{1/2}$, with $n_{\tau}^{*}$ denoting the number of testing time frames. L(95%) and P(95%) are computed similarly to those in (21) - (22), but with the predictive interval of $z$ replaced by that of $y$, where $y$ represents the velocity in each direction.